# ThinkVideo: High-Quality Reasoning Video Segmentation with Chain of Thoughts

### v2

@@ 4.3 Ablation Study
Analysis of hyper-parameter $\xi$. We sample a keyframe candidate for every $\xi$ frames, then concatenate the keyframe candidates into a grid image for the GPT-4o CoT process. In Tab. 7, we compare the reasoning VOS results on the ReasonVOS benchmark [4] with different $\xi$ values. Our empirical results showcase that GPT-4o achieves comparable performance when $T^{\prime}\leq 8$ (i.e., $\xi\geq\lfloor\frac{T-1}{8}\rfloor+1$), while the image reasoning capability drastically decreases with $\xi=\lfloor\frac{T-1}{16}\rfloor+1$. Therefore, to preserve the temporal information in the video without sacrificing GPT-4o reasoning accuracy, we recommend selecting a $\xi$ value such that $4\leq T^{\prime}\leq 8$.
Analysis of CoT process. We examine the CoT process of the MLLM keyframe selector. Specifically, we show in Fig. 7 that the CoT process significantly improves the reasoning capability. In Reasoning VIS, thinking without CoT may result in missing some instances that require more effort to find. In Online Reasoning VOS, thinking without CoT degrades the performance of the keyframe justification, possibly leading to wrong classification of binary selectivity. All of these experiments point to the soundness of our proposed CoT reasoning.
Choice of online keyframe selector. We evaluate the performance of the online ThinkVideo with different keyframe selectors $\mathcal{F}_{key}$. We use the Refer-DAVIS-17 [58] benchmark for this ablation study. Tab. 7 reports the performance and average time cost of the CoT process, where LLaVA1.5-7B [40], Qwen-2.5-VL-3B [3], and GPT-4o [29] achieve comparable quantitative results. Therefore, we choose to adopt LLaVA1.5-7B in the online ThinkVideo framework.
Limitations. While ThinkVideo demonstrates strong performance in reasoning video segmentation through zero-shot Chain-of-Thought prompting, a key limitation of ThinkVideo is its reliance on pre-trained MLLMs without task-specific fine-tuning. As a result, its performance is bounded by the general reasoning capabilities of these models. We believe that fine-tuning MLLMs for temporally-aware multimodal reasoning could further improve segmentation accuracy.
Figure 6: Visual results of the original Reasoning VIS and its online version for Reasoning VOS. The original ThinkVideo can be used for instance-level reasoning video segmentation, while the online extension is preferred when the object of interest changes during the tracking process. Zoom in for details and refer to the appendix for more examples.
@@ 5 Conclusion
In this paper, we presented ThinkVideo, a novel zero-shot, training-free framework that leverages Chain-of-Thought (CoT) capabilities of multimodal large language models (MLLMs) to enhance reasoning in video segmentation tasks. Our extensive experiments on various benchmark datasets demonstrate that ThinkVideo significantly outperforms state-of-the-art methods, particularly in challenging scenarios involving complex, implicit and time sensitive queries, occlusions, and rapid object movements. Additionally, we extended the framework for online video streams, showcasing its adaptability for a broader audience. Overall, ThinkVideo sets a new benchmark in reasoning video segmentation and opens avenues for future research in zero-shot reasoning and its applications in dynamic visual environments.
Appendix A ThinkVideo CoT Prompt Details
One of the key components of the ThinkVideo framework is the chain-of-thought keyframe selection process using GPT-4o and Gemma3 (Eq. 3). To prompt GPT-4o [29] or Gemma3-12B [65], we first resize and concatenate all the keyframe candidates $\{I^{c}_{i}\}_{i=1}^{T^{\prime}}$ into a merged image $I^{c}_{merged}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3}$, where the concatenation is done on the width dimension if $W\leq H$ or on the height dimension if $H<W$, and $\max(H^{\prime},W^{\prime})=P$ for some hyperparameter $P\in\mathbb{Z}$. Then we use the following text prompt: {adjustwidth}2em2em
@@ Appendix A ThinkVideo CoT Prompt Details
“Output list: [{object__\__index: 1, keyframe: k__\__1, object__\__description: <description of the object 1 in keyframe k__\__1>}, {object__\__index: 2, keyframe: k__\__2, object__\__description: <description of the object 2 in keyframe k__\__2>}, …]”
, where each element in the list is a dictionary with three items, with object index, keyframe index, object description. k is the k-th keyframe in the grid image. object__\__index is a numbering integer starting with 1. object__\__description implies the description for that object in a particular frame, helping the model to find the object in that particular frame. For example, a valid element in an output list can be like ‘Output list: [{object__\__index: 1, keyframe: 4, object__\__description: “the man at the top left corner of the image”}]’. You have to include all objects that fulfill the requirements in the user query. Include the objects even if it is only partially visible. While choosing the keyframe for any object, you should prioritize those frames where objects are not overlapped. This will help model to better recognize the object. Keep the output list in text format. Don’t use json formatting. The output list begins with the prefix “Output list: ”, followed by a square bracket with multiple curly brackets. The square bracket should be in the same line, following the format “Output list: […]”. Don’t start with a new line.
Here is a grid image with {num__\__keyframes} keyframes. The user query is “{query}”. Follow the instruction and output the index of the best keyframe.’
[…]
Appendix B ThinkVideo (Online) CoT Prompt Details
In the online version of ThinkVideo, an incoming frame $I_{t}$ will be sent to the MLLM keyframe selector (i.e., LLaVA-7B [40]) for binary selectivity when $t=n\xi+1$. Below is the CoT prompt for our online ThinkVideo: {adjustwidth}2em2em
@@ Appendix C Hardware Details
ThinkVideo with GPT-4o is implemented with OpenAI API and one RTX-4090 GPU. Other experiments (i.e., the ThinkVideo-Gemma3-12B and the online ThinkVideo-LLaVA1.5-7B) are implemented with two RTX-4090 GPUs.
@@ Appendix D T-ReasonVOS Dataset
ReasonVOS contains 91 videos (18 from MeViS [20], 26 from MOSE [21], 37 from BURST  [2], and 10 from VIPSeg [50]) and 458 queries, which was proposed to effectively evaluate the performance of reasoning segmentation in videos [4]. While being widely used, this benchmark falls short in giving a fair evaluation on temporal reasoning ability. In particular, some considerable number of queries are temporally insensitive and thus less challenging, where the context only refers to the appearance or function of the relevant video objects unrelated to the actions or object movements. Thus, we believe a new dataset is both timely and essential in clearly demonstrating the efficacy of ThinkVideo on temporal reasoning. Specifically, we sampled 180 video-query pairs that correspond to 45 videos (8 from MeViS, 18 from MOSE, 16 from BURST, and 3 from VIPSeg) from ReasonVOS to produce a new benchmark called T-ReasonVOS. We remove time-insensitive queries with target objects visible from the beginning to the end of the video, such as “The mode of transportation capable of transporting the largest group of people”, where users can make decisions based on independent frames with image segmentation models. Surviving data in T-ReasonVOS after removing time-insensitive counterparts focus on videos where the target objects are visible only in a short interval, or queries specifying actions and object movements, e.g., “The thing that has made a 180-degree turn.” This requires models to “think” in the temporal dimension to find the correct objects of interest. The T-ReasonVOS dataset and the selection code will be released.
@@ Appendix E CoT Qualitative Results
ThinkVideo (Original).
Refer to Appendix A for the input format. Below is two qualitative result with $T^{\prime}=7$ and $T^{\prime}=4$, respectively.
@@
ThinkVideo (Online).
Refer to Appendix B for the input format. Below is the qualitative results of two sampled frames $I_{0}$ and $I_{56}$ from the same video, where $\xi$ is set to 4:
@@
  ThinkVideo Online Example 1 (LLaVA1.5-7B):
@@
  ThinkVideo Online Example 2 (LLaVA1.5-7B):
@@
Appendix F Failure Case
@@
We demonstrate a failure case of ThinkVideo in Fig. A5. Specifically, our empirical experience shows that: when multiple object instances are visible in the video, the GPT-4o fails to identify the individual instance. Instead, similar objects are packaged within an element of the output list, e.g., “the six geese near the center of the image”. This results in the further failure of the reasoning image segmentation model. Furthermore, it is difficult to separate different instances with textual descriptions, which will be sent to the image segmentation models for each instance. Eventually, only one instance is presented in the output mask sequence in this case.
Figure A5: A failure case of ThinkVideo.
Appendix G Discussion
@@
Keyframe selector choice. We adopt GPT-4o as the keyframe selector in the original ThinkVideo, while we also conduct experiments on additional open-sourced models for better reproducibility. We notice that smaller models, like LLaVA1.5-13B [40] or Qwen2.5-VL-3B [3], failed to follow our detailed CoT prompts. Gemma3-12B produces reasonable results and performs well in the Referring VOS datasets, while falling short in difficult reasoning tasks. More comprehensive experiments of open-sourced models on larger models, such as Llama4, remain unexplored due to the limited computational resources.
@@
Dataset quality. We demonstrate an example from the ReasonVOS benchmark [4] of ThinkVideo in Fig. A6. Specifically, we notice that ThinkVideo segments the hoodie when asked for an object to alleviate the cold, while the ground-truth mask sequence refers to a blanket. This showcases that some of the reasoning queries are too ambiguous even for humans. While the benchmark is not our major focus and contribution, we have attempted to avoid such ambiguous queries when sampling the T-ReasonVOS dataset.
@@
Future directions. In this work, we are excited to have made significant strides in addressing the challenges of Reasoning Video Object Segmentation (Reasoning VOS) and pushing the boundaries further into Reasoning Video Instance Segmentation (Reasoning VIS) and Online Reasoning VOS. The qualitative and quantitative results of ThinkVideo are indeed impressive, showcasing its potential to enhance segmentation tasks. However, we acknowledge that our evaluation methods for Online Reasoning VIS and Reasoning VOS were not as rigorous as desired, where we only presented the qualitative results instead of quantitative results, primarily due to the limitations of available benchmarks in these tasks. This presents an opportunity for future research, where we aim to develop more comprehensive evaluation frameworks and datasets to thoroughly assess and refine the capabilities of ThinkVideo, ensuring it continues to advance the field of video segmentation.
Social impacts. Potential misuse of the ThinkVideo application could include its use in unauthorized surveillance or the creation of deceptive media. For example, individuals or organizations could leverage the technology to track people without their consent, leading to privacy violations and potential harassment. We believe that licensing can be a useful tool to prevent such issues, where we will establish specific terms and conditions governing the use of the application by the time we release the code.
Appendix H More Comparison
Fig. A7 contains three examples from T-ReasonVOS. We show that our proposed ThinkVideo consistently outperforms the state-of-the-art VideoLISA [4] on temporally sensitive queries.

### v4

@@ 4.3 Ablation Study
Analysis of hyper-parameter $\xi$. We sample a keyframe candidate for every $\xi$ frames, then concatenate the keyframe candidates into a grid image for the GPT-4o CoT process. In Tab.˜6, we compare the reasoning VOS results on the ReasonVOS benchmark (Bai et al., 2024) with different $\xi$ values. Our empirical results showcase that GPT-4o achieves comparable performance when $T^{\prime}\leq 8$ (i.e., $\xi\geq\lfloor\frac{T-1}{8}\rfloor+1$), while the image reasoning capability drastically decreases with $\xi=\lfloor\frac{T-1}{16}\rfloor+1$. Therefore, to preserve the temporal information in the video without sacrificing GPT-4o reasoning accuracy, we recommend selecting a $\xi$ value such that $4\leq T^{\prime}\leq 8$.
@@
Analysis of CoT process. We examine the CoT process of the MLLM keyframe selector. Specifically, we show in Fig.˜8 that the CoT process significantly improves the reasoning capability. In Reasoning VIS, thinking without CoT may result in missing some instances that require more effort to find. In Online Reasoning VOS, thinking without CoT degrades the performance of the keyframe justification, possibly leading to wrong classification of binary selectivity. All of these experiments point to the soundness of our proposed CoT reasoning.
Limitations. While CoT-RVS demonstrates strong performance in reasoning video segmentation through zero-shot Chain-of-Thought prompting, a key limitation of CoT-RVS is its reliance on pre-trained MLLMs without task-specific fine-tuning. As a result, its performance is bounded by the general reasoning capabilities of these models. We believe that fine-tuning MLLMs for temporally-aware multimodal reasoning could further improve segmentation accuracy.
@@ 5 Conclusion
In this paper, we presented CoT-RVS, a novel zero-shot, training-free framework that leverages Chain-of-Thought (CoT) capabilities of multimodal large language models (MLLMs) to enhance reasoning in video segmentation tasks. Our extensive experiments on various benchmark datasets demonstrate that CoT-RVS significantly outperforms state-of-the-art methods, particularly in challenging scenarios involving complex, implicit and time sensitive queries, occlusions, and rapid object movements. Additionally, we extended the framework for online video streams, showcasing its adaptability for a broader audience. Overall, CoT-RVS sets a new benchmark in reasoning video segmentation and opens avenues for future research in zero-shot reasoning and its applications in dynamic visual environments.
Appendix A CoT-RVS CoT Prompt Details
One of the key components of the CoT-RVS framework is the chain-of-thought keyframe selection process with temporal-semantic analysis using GPT-4o and Gemma3 (Eq.˜3). To prompt GPT-4o Hurst et al. (2024) or Gemma3-12B Team et al. (2025), we first resize and concatenate all the keyframe candidates $\{I^{c}_{i}\}_{i=1}^{T^{\prime}}$ into a merged image $I^{c}_{merged}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times 3}$, where the concatenation is done on the width dimension if $W\leq H$ or on the height dimension if $H<W$, and $\max(H^{\prime},W^{\prime})=P$ for some hyperparameter $P\in\mathbb{Z}$. Then we use the following text prompt:
@@ Appendix A CoT-RVS CoT Prompt Details
“Output list: [{object_\_index: 1, keyframe: k_\_1, object_\_description: <description of the object 1 in keyframe k_\_1>}, {object_\_index: 2, keyframe: k_\_2, object_\_description: <description of the object 2 in keyframe k_\_2>}, …]”
, where each element in the list is a dictionary with three items, with object index, keyframe index, object description. k is the k-th keyframe in the grid image. object_\_index is a numbering integer starting with 1. object_\_description implies the description for that object in a particular frame, helping the model to find the object in that particular frame. For example, a valid element in an output list can be like ‘Output list: [{object_\_index: 1, keyframe: 4, object_\_description: “the man at the top left corner of the image”}]’. You have to include all objects that fulfill the requirements in the user query. Include the objects even if it is only partially visible. While choosing the keyframe for any object, you should prioritize those frames where objects are not overlapped. This will help model to better recognize the object. Keep the output list in text format. Don’t use json formatting. The output list begins with the prefix “Output list: ”, followed by a square bracket with multiple curly brackets. The square bracket should be in the same line, following the format “Output list: […]”. Don’t start with a new line.
Here is a grid image with {num_\_keyframes} keyframes. The user query is “{query}”. Follow the instruction and output the index of the best keyframe.’
where $\{\textit{num\_keyframes}\}$ and $\{\textit{query}\}$ respectively refer to the number of keyframe candidates $T^{\prime}$ and user query $q$ in Sec.˜3.2. Lastly, we parse the output list to retrieve the keyframe selectivity $S$.
Appendix B CoT-RVS (Online) CoT Prompt Details
In the online version of CoT-RVS, an incoming frame $I_{t}$ will be sent to the MLLM keyframe selector (i.e., LLaVA-7B Liu et al. (2023a)) for binary selectivity when $t=n\xi+1$. Below is the CoT prompt for our online CoT-RVS:
@@ Appendix C Hardware Details
CoT-RVS with GPT-4o is implemented with OpenAI API and one RTX-4090 GPU. Other experiments (i.e., the CoT-RVS-Gemma3-12B and the online CoT-RVS-LLaVA1.5-7B) are implemented with two RTX-4090 GPUs.
@@ Appendix D T-ReasonVOS Dataset
ReasonVOS contains 91 videos (18 from MeViS Ding et al. (2023a), 26 from MOSE Ding et al. (2023b), 37 from BURST  Athar et al. (2023), and 10 from VIPSeg Miao et al. (2022)) and 458 queries, which was proposed to effectively evaluate the performance of reasoning segmentation in videos Bai et al. (2024). While being widely used, this benchmark falls short in giving a fair evaluation on temporal reasoning ability. In particular, some considerable number of queries are temporally insensitive and thus less challenging, where the context only refers to the appearance or function of the relevant video objects unrelated to the actions or object movements. Thus, we believe a new dataset is both timely and essential in clearly demonstrating the efficacy of CoT-RVS on temporal reasoning. Specifically, we sampled 180 video-query pairs that correspond to 45 videos (8 from MeViS, 18 from MOSE, 16 from BURST, and 3 from VIPSeg) from ReasonVOS to produce a new benchmark called T-ReasonVOS. We remove time-insensitive queries with target objects visible from the beginning to the end of the video, such as “The mode of transportation capable of transporting the largest group of people”, where users can make decisions based on independent frames with image segmentation models. Surviving data in T-ReasonVOS after removing time-insensitive counterparts focus on videos where the target objects are visible only in a short interval, or queries specifying actions and object movements, e.g., “The thing that has made a 180-degree turn.” This requires models to “think” in the temporal dimension to find the correct objects of interest. The T-ReasonVOS dataset and the selection code will be released.
@@ Appendix E CoT Qualitative Results
CoT-RVS (Original).
Refer to Appx.˜A for the input format. Below is two qualitative result with $T^{\prime}=7$ and $T^{\prime}=4$, respectively.
@@
CoT-RVS (Online).
Refer to Appx.˜B for the input format. Below is the qualitative results of two sampled frames $I_{0}$ and $I_{56}$ from the same video, where $\xi$ is set to 4:
@@
CoT-RVS Online Example 1 (LLaVA1.5-7B):
@@
CoT-RVS Online Example 2 (LLaVA1.5-7B):
@@
Appendix F Efficiency and Cost Analysis
While efficiency is not our main focus, as CoT’s complex reasoning usually requires higher computational complexity, we include the efficiency and cost analysis with each CoT-RVS component and the comparison with VideoLISA for your reference. All latency and memory measurements are based on 324-sized frames and 8 keyframe candidates for the MLLM selector. The API cost only applies to GPT-4o, which is a closed-source model. Results are reported on RTX4090.
In specific, Tab.˜A3 includes the latency, local memory footprint, and API cost of each module used in CoT-RVS and its online version. The end-to-end analysis of CoT-RVS is further presented in Tab.˜A3.
For the online CoT-RVS, the per-frame latency within a $\xi$-frame sampling window is approximated by $\frac{\alpha+S\beta+(\xi-1)\gamma}{\xi}$, where $\alpha$ is the MLLM inference latency, $\beta$ is the segmentation agent latency, $\gamma$ is the tracking latency per frame, and $S$ is the binary indicator for frame selection. This yields an upper bound of $\frac{\alpha+S\beta}{\xi}+\gamma$ for average latency, where $\gamma$ is relatively small ($0.04$ in Tab.˜A3) and $\xi$ is a hyper-parameter. Please refer to Tab.˜A3 for actual latency measurements across different MLLMs.
Table A1: Latency of each CoT-RVS component
Module Latency(s) Local Memory footprint (GB) API cost (US\$)
LLaVA1.5-7B 1.71 15.6 -
Gemma3-12B 63.95 24.8 -
GPT-4o 12.01 - 0.012
Seg-Zero 4.57 18.1 -
SAM2 0.04 2.6 -
Average per-frame tracking latency.
Table A2: End-to-end cost comparison with VideoLISA.
Method Latency(s) Local Memory footprint (GB) API cost (US\$)
VideoLISA (OTSA) 3.46 11.7 -
CoT-RVS-Gemma3-12B 69.79 24.8 -
CoT-RVS-GPT-4o 17.85 18.1 0.012
Table A3: Ablation study for online keyframe selector on Refer-DAVIS-17. * varies depending on the OpenAI server busyness. is implemented on an RTX4090.
Model $\mathcal{J\&F}$ Avg. Time (sec.)
Qwen-2.5-VL-3B Bai et al. (2025) 72.5 0.18
LLaVA1.5-7B Liu et al. (2023a) 73.9 0.24
GPT-4o Hurst et al. (2024) 73.2 0.27*
Appendix G Robustness Analysis
The quality of keyframe selection is influenced by the choice of MLLM. To evaluate the adherence of MLLM to our CoT prompt in offline setting, we compute the percentage of outputs where the MLLM produces at least one valid object instance per expression (termed "valid outputs"). Results in Tab.˜A4 show that larger, closed-source MLLMs such as GPT-4o offer significantly more robust and reliable reasoning, while smaller models like LLaVA1.5-7B and Qwen2.5-VL-3B fail to follow the output format as described in the CoT prompt (refer to Appx.˜A), thereby resulting in errors of data parsing. Prior works often underutilize this potential of large closed-source models in video segmentation tasks.
Table A4: Robustness analysis of CoT-RVS (original) with different keyframe selector models.
Dataset MLLM #Valid outputs/#Expressions Percentage $\mathcal{J\&F}$ $\mathcal{J}$ $\mathcal{F}$
MeViS (val_u) LLaVA1.5-7B 0/785 0% - - -
Qwen2.5-VL-3B 0/785 0% - - -
Gemma3-12B 715/785 91.1% 52.6 48.2 57.0
GPT-4o 761/785 96.8% 60.1 56.0 64.2
ReasonVOS LLaVA1.5-7B 0/785 0% - - -
Qwen2.5-VL-3B 0/785 0% - - -
Gemma3-12B 273/458 59.6% 50.7 47.5 54.0
GPT-4o 444/458 96.9% 65.5 62.4 68.7
Incorrect output format.
Appendix H Modularity Analysis
We have demonstrated in main paper that GPT-4o and Gemma3 can be used as MLLM keyframe selector. In this appendix, we further validate the modularity of CoT-RVS, where we adopt LISA Lai et al. (2023) and Cutie Cheng et al. (2024) alternatives to $F_{seg}$ and $F_{vid}$, respectively. We conduct this ablation under the Refer-DAVIS-2017 benchmark with Gemma3-12B as keyframe selector. Tab.˜A5 shows that both Seg-Zero and SAM2 in our default CoT-RVS can be substituted without significantly sacrificing the performance.
Table A5: Ablation study on Ref-DAVIS-17 using CoT-RVS-Gemma3-12B with different $F_{seg}$ and $F_{vid}$.
$F_{seg}$ $F_{vid}$ $\mathcal{J\&F}$ $\mathcal{J}$ $\mathcal{F}$
LISA-13B-Llama2 Cutie 68.8 65.0 72.5
LISA-13B-Llama2 SAM2 71.1 67.3 75.0
Seg-Zero-7B Cutie 72.1 68.5 75.7
Seg-Zero-7B SAM2 74.6 70.9 78.3
Appendix I Failure Case
@@
We demonstrate a failure case of CoT-RVS in Fig.˜A5. Specifically, our empirical experience shows that: when multiple object instances are visible in the video, the GPT-4o fails to identify the individual instance. Instead, similar objects are packaged within an element of the output list, e.g., “the six geese near the center of the image”. This results in the further failure of the reasoning image segmentation model. Furthermore, it is difficult to separate different instances with textual descriptions, which will be sent to the image segmentation models for each instance. Eventually, only one instance is presented in the output mask sequence in this case.
Figure A5: A failure case of CoT-RVS.
Appendix J Discussion
@@
Keyframe selector choice. We adopt GPT-4o as the keyframe selector in the original CoT-RVS, while we also conduct experiments on additional open-sourced models for better reproducibility. We notice that smaller models, like LLaVA1.5-13B Liu et al. (2023a) or Qwen2.5-VL-3B Bai et al. (2025), failed to follow our detailed CoT prompts. Gemma3-12B produces reasonable results and performs well in the Referring VOS datasets, while falling short in difficult reasoning tasks. More comprehensive experiments of open-sourced models on larger models, such as Llama4, remain unexplored due to the limited computational resources.
@@
Dataset quality. We demonstrate an example from the ReasonVOS benchmark Bai et al. (2024) of CoT-RVS in Fig.˜A6. Specifically, we notice that CoT-RVS segments the hoodie when asked for an object to alleviate the cold, while the ground-truth mask sequence refers to a blanket. This showcases that some of the reasoning queries are too ambiguous even for humans. While the benchmark is not our major focus and contribution, we have attempted to avoid such ambiguous queries when sampling the T-ReasonVOS dataset.
@@
Future directions. In this work, we are excited to have made significant strides in addressing the challenges of Reasoning Video Object Segmentation (Reasoning VOS) and pushing the boundaries further into Reasoning Video Instance Segmentation (Reasoning VIS) and Online Reasoning VOS. The qualitative and quantitative results of CoT-RVS are indeed impressive, showcasing its potential to enhance segmentation tasks. However, we acknowledge that our evaluation methods for Online Reasoning VIS and Reasoning VOS were not as rigorous as desired, where we only presented the qualitative results instead of quantitative results, primarily due to the limitations of available benchmarks in these tasks. This presents an opportunity for future research, where we aim to develop more comprehensive evaluation frameworks and datasets to thoroughly assess and refine the capabilities of CoT-RVS, ensuring it continues to advance the field of video segmentation.
Social impacts. Potential misuse of the CoT-RVS application could include its use in unauthorized surveillance or the creation of deceptive media. For example, individuals or organizations could leverage the technology to track people without their consent, leading to privacy violations and potential harassment. We believe that licensing can be a useful tool to prevent such issues, where we will establish specific terms and conditions governing the use of the application by the time we release the code.
Appendix K More Comparison
Fig.˜A7 contains three examples from T-ReasonVOS. We show that our proposed CoT-RVS consistently outperforms the state-of-the-art VideoLISA Bai et al. (2024) on temporally sensitive queries.